(sinA)x 2 +2(cosA)x+(cosA+sinA)½ if this is a square of linear fn FIND GENRL VALUS FR A

                                         (sinA)x2+2(cosA)x+(cosA+sinA)½ 


if this is a square of linear fn FIND GENRL VALUS FR A

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = (sinA)x^{2}+(cosA)(2x)+(cosA+sinA)^{1/2}
f(x) \rightarrow (x+1)^{2}
sinA =1...(1)
cosA =1….(2)
(cosA+sinA)^{1/2} =1…..(3)
From (1)
A = n\pi + \pi /2
From (2)
A = n\pi
From(3)
cosA+sinA = 1
(2)^{1/2}sin(A+\pi /4) = 1
sin(A+\pi /4) = \frac{1}{2^{1/2}}
A + \pi /4 = n\pi + \pi /4
A = n\pi
There is not any intersection between solution ot (1), (2) and (3)
A is an null set.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty